OpenStudy (anonymous):

The plane that contains the line x : 3t, f : 1 + t, z, :2t and is parallel to the intersection of the planes y + z : - 1 and2x -.r'- *=0.

OpenStudy (amistre64):

might need some clarification on what you are actually posting

OpenStudy (anonymous):

find the equation of the plane passing through the given points The plane that contains the line x : 3t, f : 1 + t, z, :2t and is parallel to the intersection of the planes y + z : - 1 and2x -.r'- *=0.

OpenStudy (anonymous):

sorry this was complete question i want to ask

OpenStudy (amistre64):

The plane that contains the line x = 3t f = 1 + tz, 2t im having trouble with your notation im wanting to see it as: x = 3t y = 1+t z = 2t but i just can be sure that is what its saying

OpenStudy (amistre64):

is it possible to post a picture of the question?

OpenStudy (anonymous):

wait im sending you

OpenStudy (anonymous):

find the equation of the plane that passing through the points ... further question is written in atachment

OpenStudy (amistre64):

beforehand, some thoughts we most likely have 2 lines that are skew, the direction vectors of these lines will need to be crossed, to find a normal vector that is perpendicular to them both. this is then the normal vector for the plane, and we can use the anchor point of the first line to establish the equation with

OpenStudy (amistre64):

the picture is much clearer to read thnx

OpenStudy (anonymous):

thanks should be mine

OpenStudy (amistre64):

we determine the line of crossing planes by crossing their own normals, and anchoring it to a point. but we only need the direction vector itself and the line is immaterial to us this make sense?

OpenStudy (anonymous):

sir we have parametric equations of line and two eq.s of plane

OpenStudy (amistre64):

plane1: 0x + 1y + 1z = -1 plane2: 2x - 1y + 1z = 0 what are our normal vectors for these planes?

OpenStudy (anonymous):

(0,1,1) (2,-1,1)

OpenStudy (amistre64):

good, and we need to cross them ... what is our cross?

OpenStudy (anonymous):

cross product ???

OpenStudy (amistre64):

yes, the cross product of the normal vectors

OpenStudy (anonymous):

2i +2j-2k

OpenStudy (amistre64):

very good, since length is immaterial to us: (1,1,-1) is fine. or (2,2,-2) any scaled version of it

OpenStudy (anonymous):

sir can i have your personal contact type thing ? to contact to you for study ?

OpenStudy (anonymous):

since length is immaterial to us: (1,1,-1) is fine. or (2,2,-2) any scaled version of it i could not understand sir ?

OpenStudy (amistre64):

now we have 2 lines to play with L1 L2 x = 0 + 3t x = a + 1t y = 1 + 1t y = b + 1t z = 0 + 2t z = c - 1t direction (3,1,2) direction (1,1,-1) we need the cross product of the direction vectors now

OpenStudy (amistre64):

we want direction, not magnitude

OpenStudy (amistre64):

the vector (2,2,-2) is equal to 2(1,1,-1)

OpenStudy (amistre64):

use either one ... both are the same direction

OpenStudy (anonymous):

-3i-5j+2k cross product

OpenStudy (amistre64):

good, (-3,5,2) is the normal to our desired plane using the anchor point from line 1 (L1) we can create the equation of the plane what was our anchor point?

OpenStudy (anonymous):

P(0,1,0)

OpenStudy (amistre64):

good, L1 is passing thru the point (0,1,0)  we have a point and a normal for our plane now